rssLink RSS for all categories
 
icon_red
icon_green
icon_orange
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_green
icon_orange
icon_orange
icon_orange
icon_green
icon_blue
icon_red
icon_orange
icon_green
icon_blue
icon_green
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_red
 

FS#3309 — Roubaix/Frankfurt

Attached to Project— Network and racks
Incident
Entire OVH Network
CLOSED
100%
Le lien est down entre Roubaix et Frankfurt.
On cherche l'origine du probleme.
Date:  Monday, 17 August 2009, 20:05PM
Reason for closing:  Done
Comment by OVH - Monday, 17 August 2009, 11:34AM

Le trafic est reparti sur Amsix, Linx au lieu de Decix.
Zurick, Milano, Prague, Vienne passent par Lyon/Marseille.


Comment by OVH - Monday, 17 August 2009, 12:29PM

La fibre est cassée après Bruxelles vers Liege.

Field engineers are currently on site and are currently performing tests.


Comment by OVH - Monday, 17 August 2009, 13:24PM

Measurements show fibre cut is located 68,5 km from Frankfurt. Engineers are now on route to perform OTDR to
provide a more accurate location for the fibre team. ETA to fibre break location is 12:00 UTC/ 14:00 CEST.


Comment by OVH - Monday, 17 August 2009, 15:06PM

The break has been confirmed due to civil works. Digging has taken place for the installation of new water pipes. The preparations
for splicing have begun and the current ETR is 16:00 UTC/ 18:00 CEST.


Comment by OVH - Monday, 17 August 2009, 16:58PM

Splicing has been delayed (NOC still trying to find out why) and so the ETR has slipped.
Splicing will start in around 15 minutes. New ETR to be obtained as soon as possible.


Comment by OVH - Monday, 17 August 2009, 18:40PM

up.


Comment by OVH - Monday, 17 August 2009, 18:47PM

Engineers inform first tube has been completely respliced. Second tube resplicing continues. Services that are up already will
remain stable.


Comment by OVH - Monday, 17 August 2009, 20:05PM

All services have been restored. Last part of resplicing was completed at 17:20 UCT/ 19:20
CEST. There will be no further updates to follow.